# Overdense fireworks in GOODS-N: Unveiling a record number of massive dusty star forming galaxies at z5.2 with the N2CLS

### v1

@@ 8 Accelerated galaxy evolution in the overdense structure
Chiang et al. (2017) outlined three evolutionary stages of cluster formation: an early inside-out growth phase from to , an extended period of vigorous star formation from to , and a final stage of rapid mass assembly and quenching from to . Our results are consistent with the second phase of this sequence, during which intense star formation is expected to occur across 10–20 cMpc-scale structures, contributing significantly to the stellar mass of present-day clusters. Interestingly, Lim et al. (2024), using the FLAMINGO simulation suite, suggest that the suppression of star formation in protocluster environments may already begin by , indicating that this epoch could represent a transitional phase in cluster evolution.
@@
Recent results from the FOREVER22 project (Morokuma-Matsui et al. 2025) indicate that the elevated star formation activity observed in protocluster galaxies at z6, compared to field galaxies, is primarily driven by a higher abundance of massive galaxies rather than enhanced specific star formation rates. This trend is likely a consequence of the denser environments in protoclusters, where increased galaxy interactions and sustained gas accretion promote the rapid assembly of massive systems. These massive galaxies tend to exhibit stronger dust attenuation, resulting in redder UV continuum slopes. This is consistent with the findings of Helton et al. (2024b), who reported that the UV-brightest and UV-reddest galaxies are preferentially located in denser environments with a higher number of neighboring galaxies. Together, these observations provide compelling evidence that galaxy evolution is accelerated within high-redshift overdense structures and may help explain the record number of bright DSFGs we identified in this particular structure.
@@
In the FoREVER22 project, protocluster galaxies show a combination of stellar masses and star formation rates that are somewhat lower, but still comparable to those of the DSFGs identified in our study. This places our DSFGs at the upper end of the FoREVER22 population in terms of total star formation activity. In contrast, the FLAMINGO simulations from Lim et al. (2024) produce galaxies with high star formation rates that are consistent with the SFRs measured for our sample. Furthermore, Chiang et al. (2017) estimate that the total SFR in protoclusters at is typically below 1000 M⊙ yr-1 which is about four times lower than the combined SFR of our five spectroscopically confirmed DSFGs. This unusually high level of star formation highlights the exceptional nature of this structure.
@@
In the context of the SIDES simulation (Béthermin et al. 2017; Gkogkou et al. 2023), where no environmental effects are considered, we can use this dataset to address whether the enhanced dusty star formation observed here is driven by clustering or by the presence of large gas masses. To explore this, we selected SIDES galaxies in the 2-deg2 simulation in the redshift range and with a NIKA2 1.2 mm flux greater than 0.7 mJy, which mimics the selection criteria for our sample (B25). The SFR–M⋆ relation for a NIKA2-like population is shown in Fig. 2. With the exception of N2GN_1_61, SIDES galaxies generally exhibit lower SFR and M⋆.
In SIDES, the SFR is capped at a hard limit of 1000 M⊙/yr, and each galaxy’s SFR is redrawn until it falls below this threshold.
This is explaining the lack of SFR1000 M⊙/yr galaxies in SIDES.
However, we can also notice that SIDES does not reach the high stellar masses. The most massive galaxy in SIDES has a stellar mass of log(M⋆) = 10.92. Even when considering the larger 117-deg2 SIDES-UCHUU simulation (Gkogkou et al. 2023), the most massive galaxy reaches log(M⋆) = 11.29.
SIDES employs abundance matching to obtain the stellar mass, which connects halo to galaxy stellar mass via mass functions. This approach may hinder the existence of high-stellar mass objects if they are not represented in the observed stellar mass function used to build the SIDES simulation.
Indeed, using the stellar mass function from SIDES, we expect 0.4 galaxies with log(M⋆/M10.7 at in the GOODS-N area (160 arcmin2), but we observe nine.
@@
Based on their stellar masses, the five galaxies are each expected to reside in massive dark matter halos, with a combined halo mass of M⊙, as inferred from the empirical – relation at (Behroozi et al. 2019). While it is uncertain whether all these galaxies will eventually merge into a single structure by , such an outcome would imply a descendant halo more massive than that of a typical Coma-like cluster progenitor at the same redshift, which simulations suggest is closer to M⊙ (Chiang et al. 2013). This highlights the exceptional mass scale of the environment traced by these galaxies. However, it is important to note that total mass estimates for protoclusters are sensitive to the choice of aperture and redshift interval used for their definition, influencing such comparisons across studies (Lim et al. 2024).
@@ 9 Conclusion
Using 1.2 mm observations with NIKA2 in GOODS-N, we identify ten DSFGs at , many of which are extremely faint or even undetected in deep HST and some JWST imaging. These galaxies exhibit remarkably high SFRs, substantial dust content, and significant stellar masses, placing them among the most extreme star-forming systems known at this epoch. Six are securely associated with the exceptional overdensity reported in Sun et al. (2024) and Herard-Demanche et al. (2025).
@@
We first identify a new DSFG in the overdense environment of HDF850.1 through NOEMA observations, which detect CO(4–3) and CO(5–4) lines at . In addition, another N2CLS galaxy shows a 7 emission line in FRESCO data, consistent with H at z. Combined with previously known galaxies (HDF850.1, GN10, and S3), this brings the number of spectroscopically confirmed DSFGs within the overdense structure to five. We further identify five DSFGs with photometric redshifts consistent with the overdensity.
@@
These galaxies exhibit exceptionally high star formation rates (ranging from 200 to 2000 M⊙,yr-1) and large stellar masses (log(M⋆/M⊙) = 10–12), placing most of them significantly above the star-forming main sequence at , especially under the Popesso et al. (2023) parametrization. Their properties are comparable to those of the ultra-massive galaxies analyzed by Xiao et al. (2024), suggesting an early and efficient buildup of both stellar and dust mass. They show short gas depletion timescales (with a median of 0.19 Gyr) and remarkably high stellar baryon conversion efficiencies (20%, compared to 10% locally). The observed dust masses further support the scenario of efficient dust formation. Due to their short gas depletion timescales, these DSFGs could evolve into quiescent galaxies by , assuming no new gas inflow. This scenario aligns with the early formation of massive quiescent galaxies seen in similar dense environments.
@@
Their contribution to the cosmic SFRD is significant, exceeding the total SFRD by a factor of 1.4 when considering spectroscopically confirmed galaxies, and by a factor of 4.3 when including all ten sources. These few dusty galaxies dominate the star formation within the overdensity, contributing more than the numerous Hα emitters.
@@
Our DSFGs either lie at the upper extremes of stellar mass, dust mass, and SFR, or are entirely absent from current simulations and models. This highlights the difficulty existing models face in reproducing such extreme systems. Their properties may be driven by efficient gas inflows along cosmic filaments in protocluster environments, fueling intense star formation despite short depletion timescales (e.g. Narayanan et al. 2015).
@@
We examine the spatial distribution of N2CLS galaxies within the overdensity, comparing their positions to the three filamentary structures identified by Sun et al. (2024). Four spectroscopically confirmed DSFGs lie along filament-1 and filament-2, while two others align with filament-3 and one may trace a new separate structure. The protocluster may be more extended than previously mapped, as several of our galaxies lie beyond current survey footprints. Confirming this will require spectroscopic redshifts for sources currently identified via photometric estimates. If all these galaxies merge by , they would form a halo significantly more massive than a typical Coma-like cluster progenitor at that epoch, emphasizing the exceptional mass scale of the environment traced by these galaxies.
@@
The presence of such massive, dusty, and intensely star-forming galaxies at has significant implications for our understanding of galaxy formation and evolution. It suggests that the processes leading to rapid stellar and dust mass assembly were already in place within the first billion years of cosmic history in overdense environments. Future observations, particularly with facilities capable of probing the ISM properties and kinematics of these galaxies, will be crucial in unraveling the mechanisms driving their evolution.

### v2

@@ 8 Accelerated galaxy evolution in the overdense structure
Chiang2017 outlined three evolutionary stages of cluster formation: an early inside-out growth phase from to , an extended period of vigorous star formation from to , and a final stage of rapid mass assembly and quenching from to . Our results are consistent with the second phase of this sequence, during which intense star formation is expected to occur across 10–20 cMpc-scale structures, contributing significantly to the stellar mass of present-day clusters. Interestingly, Lim2024, using the FLAMINGO simulation suite, suggest that the suppression of star formation in protocluster environments already began by , indicating that this epoch could represent a transitional phase in cluster evolution.
@@
Recent results from the FOREVER22 project (FOREVER22) indicate that the elevated star formation activity observed in protocluster galaxies at z6, compared to field galaxies, is primarily driven by a higher abundance of massive galaxies rather than enhanced sSFRs. This trend is likely a consequence of the denser environments in protoclusters, where increased galaxy interactions and sustained gas accretion promote the rapid assembly of massive systems. These massive galaxies tend to exhibit stronger dust attenuation, resulting in redder UV continuum slopes. This is consistent with the findings of Helton2024_b, who reported that the UV-brightest and UV-reddest galaxies are preferentially located in denser environments with a higher number of neighbouring galaxies. Together, these observations provide compelling evidence that galaxy evolution is accelerated within high-redshift overdense structures and may help explain the record number of bright DSFGs we identified in this particular structure.
@@
In the FoREVER22 project, protocluster galaxies show a combination of stellar masses and SFRs that are somewhat lower, but still comparable to those of the DSFGs identified in our study. This places our DSFGs at the upper end of the FoREVER22 population in terms of total star formation activity. In contrast, the FLAMINGO simulations from Lim2024 produce protocluster galaxies with high SFRs that are consistent with the SFRs measured for our sample. This is not the case for the general DSFG population modelled in FLAMINGO (kumar2025), selected based on their 850 m flux (S mJy, corresponding to the N2CLS flux cut S mJy at ), which predominantly lies on the SFMS (see Fig. 2). Furthermore, Chiang2017 estimate that the total SFR in protoclusters at is typically below 1000 M⊙ yr-1, which is about four times lower than the combined SFR of our five spectroscopically confirmed DSFGs. The unusually high level of star formation highlights the exceptional nature of this structure.
@@
In the context of the SIDES simulation (Bethermin2017; Gkogkou2023), where no environmental effects are considered, we can use this dataset to address whether the enhanced dusty star formation observed here is driven by clustering or by the presence of large gas masses. To explore this, we selected SIDES galaxies in the 2-deg2 simulation in the redshift range and with a NIKA2 1.2 mm flux greater than 0.7 mJy, which mimics the selection criteria for our sample (B25). The SFR–M⋆ relation for a NIKA2-like population is shown in Fig. 2. With the exception of N2GN_1_61, SIDES galaxies generally exhibit lower SFR and M⋆. In SIDES, the SFR is capped at a hard limit of 1000 M⊙/yr, and each galaxy’s SFR is redrawn until it falls below this threshold. This is explaining the lack of SFR1000 M⊙/yr galaxies in SIDES. However, we can also see that SIDES does not reach high stellar masses. The most massive galaxy in SIDES has a stellar mass of log(M⋆/M⊙) = 10.92. When considering the larger 117-deg2 SIDES-UCHUU simulation (Gkogkou2023), the most massive galaxy reaches log(M⋆/M⊙) = 11.29. SIDES employs abundance matching to obtain the stellar mass, which connects halo to galaxy stellar mass via mass functions. This approach may suppress the presence of high-stellar mass objects if they are not represented in the observed stellar mass function used to build the SIDES simulation. Indeed, using the stellar mass function from SIDES, we expect 0.4 galaxies with log(M⋆/M10.7 at in the GOODS-N area (160 arcmin2), whereas we observe six (or seven if N2GN_1_20 is included).
@@
Based on their stellar masses, the five galaxies are each expected to reside in massive dark matter halos, with a combined halo mass of M⊙, as inferred from the empirical – relation at (Behroozi2019). While it is uncertain whether all these galaxies will eventually merge into a single structure by , such an outcome would imply a descendant halo more massive than that of a typical Coma-like cluster progenitor at the same redshift, which simulations suggest is closer to M⊙ (Chiang2013). This highlights the exceptional mass scale of the environment traced by these galaxies. However, it is important to note that total mass estimates for protoclusters are sensitive to the choice of aperture and redshift interval used for their definition, influencing such comparisons across studies (Lim2024).
@@ 9 Conclusion
Using 1.2 mm observations of GOODS-N taken with NIKA2, we identified eight N2CLS galaxies at (five with zspec and three with zphot), many of which are extremely faint or even undetected in deep HST and some JWST imaging. These galaxies have remarkably high SFRs, substantial dust content, and significant stellar masses, placing them among the most extreme star-forming systems known at this epoch. Six are securely associated with the exceptional overdensity reported in Sun2024 and Herard-Demanche2024.
@@
We first identified a new DSFG in the overdense environment of HDF850.1 through NOEMA observations, which detect CO(4–3) and CO(5–4) lines at . In addition, another N2CLS galaxy shows a 7 emission line in FRESCO data, consistent with H at z. Combined with previously known galaxies (HDF850.1, GN10, and S3), this brings the number of spectroscopically confirmed DSFGs within the overdense structure to five. We further identified three DSFGs with photometric redshifts consistent with the overdensity.
@@
These galaxies exhibit exceptionally high SFRs (ranging from 200 to 2000 M⊙,yr-1) and high stellar masses (log(M⋆/M⊙) = 10–11), placing most of them significantly above the SFMS at , especially under the popesso2023 parametrisation. Their properties are comparable to those of the ultra-massive galaxies analysed by Xiao2024, suggesting an early and efficient buildup of both stellar and dust mass. They show short gas depletion timescales (with a median of 0.18 Gyr) and remarkably high stellar baryon conversion efficiencies (20%, compared to 10% locally). The observed dust masses further support the scenario of efficient dust formation. Due to their short gas depletion timescales, these DSFGs could evolve into quiescent galaxies by , assuming no new gas inflow. This scenario aligns with the early formation of massive quiescent galaxies seen in similar dense environments.
@@
Their contribution to the cosmic SFRD is significant: even considering that it represents only the obscured component, the five spectroscopically confirmed galaxies exceed the total SFRD from Madau2014 by a factor of 1.4, which rises to 2.6 when including all eight sources. Compared to more recent SFRD estimates at z5 (e.g. fujimoto2024), the excess is reduced to roughly unity for the spectroscopic sample and by about a factor of 2 when including photometric redshift members. These few dusty galaxies dominate the star formation budget within the overdensity, contributing more than the much larger population of H emitters.
@@
Our DSFGs either lie at the upper extremes of stellar mass, dust mass, and SFR or are entirely absent from current simulations and models. This highlights the difficulty existing models face in reproducing such extreme systems. Their properties may be driven by efficient gas inflows along cosmic filaments in protocluster environments fuelling intense star formation despite short depletion timescales (e.g. Narayanan2015; Umehata2019).
@@
We examined the spatial distribution of N2CLS galaxies within the overdensity, comparing their positions to the three filamentary structures identified by Sun2024. Four spectroscopically confirmed DSFGs lie along filament-1 and filament-2, two others align with filament-3, and two may trace new separate structures. We also identified an extremely massive (31012 M⊙) galaxy candidate at z=5.33 that lies farther away (see Appendix D). The protocluster may therefore be more extended than previously mapped, as some galaxies lie beyond current survey footprints. Confirming this will require spectroscopic redshifts for sources currently identified via photometric estimates. If all these galaxies merged by , they would form a halo significantly more massive than a typical Coma-like cluster progenitor at that epoch, emphasising the exceptional mass scale of the environment traced by these galaxies.
@@
The presence of such massive, dusty, and intensely star-forming galaxies at has significant implications for our understanding of galaxy formation and evolution. It suggests that the processes leading to rapid stellar and dust mass assembly were already in place within the first billion years of cosmic history in overdense environments. Future observations, particularly with facilities capable of probing the ISM properties and kinematics of these galaxies, will be crucial in unravelling the mechanisms driving their evolution.